OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My equation is f(x)=2(x-1)^2+21

OpenStudy (anonymous):

use the slope formula... you need 2 points average rate of change is f(b)−f(a)b−a so substitute the x values to get a corresponding f(x) then substitute into the formula

OpenStudy (anonymous):

of so looking at the equation in vertex form find F(7) F(7)=3(7−3)2+4=52 and for x = 4 F(4)=3(4−3)2+4=7 so the average rate of change is 52−77−4= same process for the standard form equation.

OpenStudy (anonymous):

oops just realised they are the same equations... so all you need to do is simplify the fraction for the average rate of change...

OpenStudy (anonymous):

you do... you actually get 2 points as shown above (7, 52) and (4, 7) so all you need to do for the average rate of change is find the slope of the interval joining the 2 points

OpenStudy (anonymous):

well you you go up to my 1st post you have (52 - 7)/(7 -4) = just evaluate it. thats the rate of change

OpenStudy (anonymous):

I am really not understanding this at all.. I'm sorry.

OpenStudy (anonymous):

you should get 15...

OpenStudy (anonymous):

15 is the answer?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thank you